OpenStudy (anonymous):

Anyone? Can a right triangle have sides with measures of 30 ft 50 ft and 60 ft? why or why not? Give an example of three side lengths that are possible and impossible for a right triangle.

OpenStudy (rea201):

Use the pythagorean theorem. 30^2+50^2=60^2 make sure it is correct

OpenStudy (anonymous):

\[a^2 + b^2 = c^2 \] \[(30)^2 + (50)^2 = (60)^2 \]

OpenStudy (anonymous):

you can test it with pythagorean theorem \[A^2+B^2 = C ^2\] 

OpenStudy (anonymous):

\[900 + 2500 = 3600\] \[3400 = 3600\] FALSE THEREFORE: YOU CAN NOT CREATE A RIGHT TRIANGLE WITH THE GIVEN SIDE LENGTHS.

OpenStudy (anonymous):

what is possible for a right triangle?

OpenStudy (anonymous):

you don't have any ideas?

OpenStudy (anonymous):

90 90 90?

OpenStudy (anonymous):

why do you think this? try it and see... 90^2 + 90^2 = 90^2

OpenStudy (anonymous):

If true, then yes you can make a right triangle. If no, then you can not make a right triangle with these side lengths.

OpenStudy (anonymous):

it cant.. what about 29^2 20^2 21^2

OpenStudy (anonymous):

It can right?

OpenStudy (anonymous):

So your side lengths would be: 29-20-21 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If you pick two numbers and square each of them, then add the squares and take the square root of the answer, you get the hypotenuse of a right triangle. a^2 + b^2 = c^2 so c = sqr rt of (a^2 + b^2 ) For example 3, 4, and 5 are lengths that satisfy this equation. And any multiple of these three should also work. Like 30, 40, and 50. Can you thing of another 3 numbers that fit the Pythagorean theorem?

OpenStudy (anonymous):

29-20-21?